LASIK eye surgery uses pulses of laser light to shave off tissue from the cornea, reshaping it. A typical LASIK laser emits a 1.0-mm-diameter laser beam with a wavelength of 193 nm. Each laser pulse lasts 10 ns and contains 1.0 m.J of light energy.

 You may want to review (Pages 484 - 485) .


 Part A

 What is the power of one laser pulse?


 Part B

 During the very brief time of the pulse, what is the intensity of the light wave?
Answer #1

P = energy / time

P = 10^-3 / (10*10^-9)

P = 1.0*10^5 W

==========

area, A = pi * r^2

A = 3.14* (0.5*10^-3)^2

A = 7.85*10^-7 m^2

Intensity I = P/ A

I = 10^5/ ( 7.85*10^-7)

I = 1.3*10^11 W/m^2

=========

do comment in case any doubt, will reply for sure.. that